OpenStudy (sugarmochi):

The roots of x^2 -6x+k=0 differ by 1. Find the value of k

OpenStudy (3mar):

May I help?

OpenStudy (3mar):

I think this would help you.

OpenStudy (eliesaab):

Let r and s be the two roots, then r+s=6 r-s=1 this gives r=7/2 an s=5/2 k = rs =35/4

OpenStudy (eliesaab):

For problems like these, try http://www.saab.org/mathdrills/act.html

OpenStudy (3mar):

@eliesaab Thank you for good explanation.